.99
How does sample size affect validity of results in a research?
A large sample reduces the variability of the estimate. The extent to which variability is reduced depends on the quality of the sample, what variable is being estimated and the underlying distribution for that variable.
The shape of the sampling distribution is always approximately normal?
welll its quiet simple to be honest 1st ASK YOUR TEACHER! that's what they are there for
1.0966
If you roll a die 60 times what is the probability you will roll a one?
You add the probabilities together. For each roll, it is 1/6; therefore, you will have 60 times 1/6, which is 60/6, which is 10. Since probability is only 1 at the most, which means that it WILL happen, I would say that the probability is 1 (or 100%).
Why is the sample mean an unbiased estimator of the population mean?
The sample mean is an unbiased estimator of the population mean because the average of all the possible sample means of size n is equal to the population mean.
Suppose probability of drawing a red marble is p.
Then p = 2*(1 - p)
that is p = 2 - 2p
or p = 2/3
So 2/3 of the 24 marbles are red
24*(2/3) = 16 red marbles.
What is the probability of obtaining 3 heads in three flips of a fair coin?
1/2 * 1/2 * 1/2 = 1/8 = 12.5%
Find the distinguishable permutations of the word Texas?
Why, sixty! (Did you know that they have taxes, in Texas?!)
The calculation is 5! = 5 X 4 X 3 X 2 X 1 = 60.
There are five choices of which letter to use in the first position, four choices for the next position, and so on.
What does P with an x through the bottom of the P mean on Christian symbols?
it is a symbol for Christianity, based of the two first letters of the word Christ in Greek. Constantine saw it in the sky before a battle and was told that by that symbol he would conquer, so he painted it on his army's shields and fought extremely well!
Short answer: 14/81
There is a 2/9 chance of picking a green marble the first time.
Since this leaves 8 marbles (only one of which is green; you took the other one), that leaves an 7/8 chance of picking any other colour the second time.
To find the overall probability, you simply multiply the two values (2/9 * 7/8 = 14/81). This is simple to understand. You can only get a green marble every 2 out of 9 tries. For each of these 2 times, you have a 7/8 chance of getting a different colour the second time. So, you have a 7/8 chance every 2/9 times. Simply multiply them to get the total number of possibilities.
You roll a number cube 150 times about how many times would you expect to roll 3?
Possible outcomes of one roll = 6
Successful outcomes = 1
Probability of success on each roll = 1/6
Expectation in 150 rolls = (1/6) x (150) = 25 times
How does a large sample size help to ensure that experimental results are reliable?
It made his actual results approach the results predicted by probability
There are 24 df.
Probability lesson represent outcomes 4th grade math?
That's not a question; indeed, it isn't even a sentence. If you're looking for a lesson to teach to 4th-graders, you might try an educators' site instead.
-3/12
3/6 or 1/2 or 50%